Sue can collect data from 5 more houses per hour than Joe. It took Sue 2 hours less to collect data from 120 houses than it took Joe to collect data from the same number of houses. What is Joe's rate of collecting data.
Answer by stanbon(75887)

Sue can collect data from 5 more houses per hour than Joe.
Joe rate = x houses/hour
Sue rate = (x+5) houses/hour
--------------------------------
It took Sue 2 hours less to collect data from 120 houses than it took Joe to collect data from the same number of houses. What is Joe's rate of collecting data.
------------
Sue time to collect 120 houses:: 120/(x+5))
-----
Joe time to collect 120 houses 120/x) = 120x
-----
120/x - 120/(x+5) = 2 hrs
-------------------------
120(x+5) - 120x = 2x(x+5)
600 = 2x^2 + 10x
x^2 + 5x - 300 = 0
------------------------
x = 15 houses per hour (Joe's rate)
x+5 = 20 houses per hour (Sue's rate)
-----------------
Cheers,
Stan H.